OpenStudy (anonymous):

Each side of a square is increasing at a rate of 4 cm/s. At what rate is the area of the square increasing when the area of the square is 81 cm2?

OpenStudy (anonymous):

related rates right?

OpenStudy (anonymous):

area of a square with side \(x\) is \(A=x^2\) differentiate with respect to time and you get \[A'=2xx'\]

OpenStudy (anonymous):

you are told \(x'=4\) and since \(A=81\) that makes \(x=\sqrt{81}=9\)

OpenStudy (anonymous):

plug them in and get \(A'\) by multiplication

OpenStudy (anonymous):

not sure if that was clear, but i hope so final answer is \[A'=2\times 9\times 4\]

OpenStudy (anonymous):

so I get my answer by multiplying 2 9 and 4

OpenStudy (anonymous):

oui

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

yw